OpenStudy (anonymous):

find sin (alpha+ beta), tan=21/20, alpha lies in quadrant 3, cos beta=(-8/17), b lies in q2 please give me details on how to do this, and what identity was used. thank you

OpenStudy (tkhunny):

Quadrant III tan(x) > 0 sin(x) < 0 cos(x) < 0 tan(x) = sin(x)/cos(x) sin(x+y) = sin(x)cos(y) + cos(x)sin(y) That's about all you need. Go! Let's see what you get.

OpenStudy (tkhunny):

Oh, you may also need \(\sin^{2}(x) + \cos^{2}(x) = 1\)

OpenStudy (anonymous):

I am still so confused :( where do I put the values for tan alpha and cos beta?

OpenStudy (tkhunny):

You just have to put in the time and think it through. All the pars are there. \(\alpha\) is in III tangent > 0 and we have \(\tan(\alpha) = 21/20\) sine < 0 cosine < 0 \(\beta\) is in II tangent < 0 sine > 0 cosine < 0 and we have \(\cos(\beta) = -8/17\) \(\sin(\beta) = \sqrt{1 - (-8/17)^{2}}\) \(\sin(\alpha)/\cos(\alpha) = \tan(\alpha) = 21/20\) \(\sin^{2}(\alpha) + \cos^{2}(\alpha) = 1\) Use these to find \(\sin(\alpha)\) and \(\cos(\alpha)\)